OpenStudy (goten77):

so (sqrt(x))^2 = abs(x) the reason is that that square part will make it a positive no matter what the inside is? also does this then mean sqrt(x^2) = x since the ^2 is not on the outside?

OpenStudy (goten77):

\[\sqrt{x}^{2} = \left| x \right|\]

OpenStudy (zzr0ck3r):

correct

OpenStudy (anonymous):

if x <0 then (sqrt(x))^2 != abs(x) as left hand side is -ve but RHS is +ve

OpenStudy (goten77):

\[\sqrt{x ^{2}} = x\]

OpenStudy (zzr0ck3r):

not always. imagine if x = -1

OpenStudy (zzr0ck3r):

so this is true when x >= 0

OpenStudy (zzr0ck3r):

things can get strange when in the world of powers example 1 = (-1)^(2/10) = (-1)^(1/5) = -1

OpenStudy (goten77):

 so when x<0 then x=x ?

OpenStudy (zzr0ck3r):

I guess it depends on if you take the square root first or wquare it first sqrt((-1)^2) = sqrt(1) = 1

OpenStudy (zzr0ck3r):

my calculator says sqrt((-1)^2) = 1 but it says (-1)^(2/2) = -1 so go figure...

OpenStudy (goten77):

well the top part is = to 1 and the bottom part is like wtf lol calculators get tricked ez

OpenStudy (zzr0ck3r):

not ti-89s explain this then 1 = (-1)^(2/10) = (-1)^(1/5) = -1 there is more to powers than we are tought in normal math class. some sort of order of operations...